What are the dimensions of a rectangular field of area A, that requires the least

amount of fencing?


1
Expert's answer
2021-12-17T04:57:58-0500

length of fencing:

"L=2(x+y)"

area:

"A=xy\implies y=A/x"


"L=2(x+A\/x)"

"L'(x)=2(1-A\/x^2)=0"


"x=y=\sqrt A"